# Diffusivity-Free Turbulence in Tabletop Rotating Rayleigh-Bénard Convection Experiments

### v1

@@
Table 2: Power law fits corresponding to the data shown in Fig. 3. Prefactors ($A$, $B$, $C$) and exponents ($\alpha$, $\beta$, $\gamma$) of unity correspond to the result of the plane layer asymptotically reduced models of (Julien et al., 2012).
@@
Measurements of the inertial oscillation-dominated range of liquid metal RRBC in closed cylindrical laboratory containers reveal that asymptotic DF heat and momentum transfer can be realized on the laboratory scale. We observe these trends in measurements of internal flow velocities and thermal fluctuations regardless of tank geometry. In contrast, global heat transfer is significantly influenced by wall modes in narrow, low aspect ratio cells, with asymptotic trends predominantly found in wide, high aspect ratio cells that better resemble plane layer geometry. If wall modes can be suppressed (cf. Terrien et al., 2023), it may be possible to achieve asymptotic heat transfer scalings in taller, more slender $\Gamma\lesssim 1$ tanks. It may be possible to further investigate DF scaling behaviors using lower $Pr$ fluids, such as liquid sodium, in which $Ra_{S}/Ra_{O}$ attain larger values. By verifying here that diffusivity-free convection exists in a standard desktop-scale RRBC experimental cell, it is now possible to accurately and confidently model the convective turbulence that exists in rapidly rotating planetary and stellar interiors.
Data and Code Availability
All data generated or analyzed during this study are included in this article (and its supplementary information). Custom code is not used in the analysis of the data.
Author Contributions
J.A.A. and J.M.A. conceived the idea for the study; Y.X. constructed this iteration of the experimental device; J.A.A. and Y.X. conducted all laboratory experiments; T.V. contributed to Doppler data collection; J.A.A. conducted the data analysis; J.A.A., S.H., K.J., and J.M.A. conducted the theoretical analysis; J.A.A. and J.M.A. wrote the original manuscript with revisions contributed by Y.X., T.V., and S.H.
Methods
Experimental device. The laboratory device used here, RoMag, has been used in a number of previous liquid metal convection studies (King and Aurnou, 2013, 2015; Aurnou et al., 2018; Vogt et al., 2018, 2021; Grannan et al., 2022; Xu et al., 2022, 2023), two of which produced results included in Fig. 2 and Fig. 3 for analysis (King and Aurnou, 2015; Vogt et al., 2021). For the new data acquired in this study, the setup consists of a right cylindrical cell of inner diameter $D=2R=19.68$ cm and height $H=40$ cm to yield an aspect ratio of $\Gamma=D/H=0.49$. The cell sidewall is made of stainless steel, the top and bottom plate boundaries are made of copper, and the working fluid is liquid gallium ($Pr\approx 0.027$). Layers of aerogel insulation (approximately 3 cm thick total) surround the cell sidewall, while a 7.62 cm thick calcium silicate insulation block sits below the lower plate boundary to mitigate heat losses. A non-inductively wound electrical heating pad at the bottom of the cell supplies a fixed flux of heating power ranging from $P=10$ to 2000 W, while the top of the cell is maintained at a fixed temperature using a recirculating chiller. The generated temperature gradients range from approximately $\Delta T=1$ to 60 K. The cell is mounted on a platform that rotates at rates ranging from $\Omega=0.105$ to $\Omega=3.67$ rad/s (1 to 35 revolutions per minute). A systematic survey was performed such that the Rayleigh and Ekman numbers each vary by approximately two orders of magnitude ($8\times 10^{6}<Ra<1\times 10^{9}$ and $3\times 10^{-7}<E<1\times 10^{-5}$).
Data acquisition. Twelve thermistors are located within the top and bottom plate boundaries to measure the global temperature gradient,
$\Delta T=\overline{T}_{bot}-\overline{T}_{top},$ (6)
which is used in the calculation of the Rayleigh and Nusselt numbers (e.g., King et al., 2012; Xu et al., 2022; Abbate and Aurnou, 2023). Six thermistors contribute to each the top and bottom boundary temperature calculation following
$\overline{T}_{top,\ bot}=\frac{1}{6N_{t}}\sum_{i=1}^{6}\sum_{j=1}^{N_{t}}T_{ij}\ \pm\ \frac{\overline{q}d}{k_{Cu}},$ (7)
where $\overline{T}_{top}$ is determined using only sensors in the top plate, $\overline{T}_{bot}$ is determined from those in the bottom, and $N_{t}$ is the length of the time series. The thermistors are embedded within the copper plates at a location $d=2$ mm away from the fluid-plate interface, so the measured temperature is adjusted to correct for conduction over length $d$ via the second term in (7), where $\overline{q}$ is the time-averaged heat flux given by (11) and $k_{Cu}$ is the thermal conductivity of the copper plate. The temperature-dependent fluid material properties are calculated using the mean temperature,
$\overline{T}=(\overline{T}_{bot}+\overline{T}_{top})/2.$ (8)
The input heat flux, $q_{in}=P_{in}/A$, is set by the power supply where $P_{in}$ is the wattage of the input heating power and $A$ is the active surface area of the fluid-plate interface. This value is confirmed with measurements of current, $I$, and voltage, $V$, at the location of the heating pad to calculate $P_{in}=IV$. Thermistors located within the layers of aerogel insulation are used to measure the heat lost through the cell sidewall. This loss is calculated as
$P_{sw,loss}=\frac{2\pi k_{ins}H(T_{sw}-T_{ins})}{\log(R_{ins}/R_{sw})},$ (9)
where $k_{ins}$ is the thermal conductivity of aerogel, $T_{sw}$ is the temperature measured at cell sidewall, $T_{ins}$ is the temperature measured after one layer of insulation, $R_{sw}$ is the radial position of the sidewall thermistor, and $R_{ins}$ is the radial position of the insulation thermistor. This loss is subtracted from the power measured at the location of the heating pad. Some power is additionally lost to vertical conduction through the stainless steel sidewall. This is estimated via
$P_{sw,cond}=\frac{\pi(R_{sw}^{2}-R^{2})k_{sw}\Delta T}{H},$ (10)
where $R$ is the inner radius of the cylinder and $k_{sw}$ is the thermal conductivity of the stainless steel sidewall. The total heat flux through the fluid is thus
$\overline{q}=\frac{\overline{P}_{in}-\overline{P}_{sw,loss}-\overline{P}_{sw,cond}}{\pi R^{2}},$ (11)
where each overlined quantity is time-averaged. This $\overline{q}$ is then used in the calculation of the Nusselt number.
Five internal thermistors are located within the fluid interior of the cell to measure bulk temperature fluctuations, $\theta$. Internal thermistor locations are shown in Fig. 1b and Supplementary Fig. 4b. Each internal thermistor measures the local temperature at its respective location over the course of the experiment. The standard deviation of the temperature time series provides the thermal fluctuation at that location (Fig. 1c). The root-mean-square (rms) value of the five thermistors’ measured fluctuations is calculated to estimate the characteristic temporally and spatially averaged temperature fluctuation, $\theta$, in the convecting fluid bulk. The equation is thus given as
$\theta=\sqrt{\frac{1}{5N_{t}}\sum_{i=1}^{5}\sum_{j=1}^{N_{t}}(T_{i,j}-\overline{T}_{i})^{2}}.$ (12)
This value is then normalized by $\Delta T$ in our analysis.
[…]
$u_{z}=\sqrt{\frac{1}{N_{z}N_{t}}\sum_{i=1}^{N_{z}}\sum_{j=1}^{N_{t}}u_{i,j}^{2}}\ ,$ (13)
where $u_{i,j}$ is each velocity measurement at space element $i$ and time element $j$. The length of the time series is represented by $N_{t}$ and the length of the spatial array is given by $N_{z}$.
Modes of convection. Convective instabilities arise in a number of different ways in rotating systems. Here, three modes of convection are relevant to the experiments conducted.
The primary mode is known as the steady or stationary mode. This type of instability occurs when thermal buoyancy overcomes the stabilizing effect of rotation, such that fluid warmed by bottom boundary rises, and fluid cooled by the top boundary sinks. This mode is therefore fueled by thermal fluxes through the boundaries. The critical value of $Ra$ for which steady convection onsets is dependent on the Ekman number following
$Ra_{S}^{\infty}=8.7E^{-4/3}$ (14)
in a horizontal plane layer in the limit of infinitely low $E$ (Chandrasekhar, 1961; Aurnou et al., 2018), or
$Ra_{S}=(8.7-9.63E^{1/6})E^{-4/3}$ (15)
for a planar system rotating at a finite rate ($E>0$) and no-slip top and bottom boundaries (Niiler and Bisshopp, 1965; Kunnen, 2021). In this work, expression (15) is employed.
In systems with lateral boundaries, convective instabilities additionally develop at the sidewall in the form of waves of alternating upward and downward flow that travel around the perimeter of the cell. These are known as wall modes. For a no-slip, thermally insulated system, wall modes are predicted to onset at a critical Rayleigh number given by
$Ra_{W}=\pi^{2}\sqrt{6\sqrt{3}}E^{-1}+46.55E^{-2/3}$ (16)
(Zhang and Liao, 2009). These modes are strongest at the periphery of the tank and aren’t expected to significantly affect the bulk interior flow quantities (i.e., $u_{z}$, $\theta$). However, they do contribute to the total global heat transport ($Nu$) for the system.
Specific to rotating fluids with $Pr<0.68$, convective instability can also develop in the form of thermally driven inertial oscillations known as oscillatory modes (Zhang and Roberts, 1997). This instability onsets at a lower $Ra$ than the critical value for steady modes. The prediction for onset in a horizontally infinite plane layer follows
$Ra_{O}^{\infty}=17.4\left(\frac{E}{Pr}\right)^{-4/3}$ (17)
(Chandrasekhar, 1961). Oscillatory onset predictions in finite cylindrical geometry subject to non-slip experimental boundary conditions, $Ra_{O}^{cyl}$, are given in Zhang and Liao (2009), verified through comparisons with simulations at $E=10^{-4}$, $\Gamma=2$ and $10^{-6}\leq Pr\leq 0.05$. In contrast to steady and wall modes, which are directly linked to the system boundaries, oscillatory modes are driven by the internal temperature gradient, and are therefore independent of the boundaries. Isolating the oscillatory mode can uniquely enable convection that is free from diffusive boundary effects.
Taking the ratio of the critical onset prediction for the stationary mode to that of the oscillatory mode gives
$\frac{Ra_{S}^{\infty}}{Ra_{O}^{\infty}}=\frac{1}{2}Pr^{-4/3}.$ (18)
Equation (18) shows that the smaller the Prandtl number is, the larger the separation between the onset of each mode. Liquid gallium, at $Pr\simeq 0.027$, yields $Ra_{S}^{\infty}/Ra_{O}^{\infty}\simeq 60$, so we can access oscillatory-dominated convection for approximately 1-2 orders of magnitude in the Rayleigh number. Fluids with smaller Prandtl numbers, such as liquid sodium with $Pr\simeq 0.001$, would provide an even larger range in which oscillatory convection can be studied with $Ra_{S}^{\infty}/Ra_{O}^{\infty}\simeq 10^{3}$, yielding approximately 3 orders of magnitude in $Ra$ before steady convection onsets.
The ratio of the wall mode onset prediction to that of the oscillatory mode simplifies to
$\frac{Ra_{W}}{Ra_{O}^{\infty}}=(1.83E^{1/3}+2.74E^{2/3})\ Pr^{-4/3}.$ (19)
According to (19), the Ekman number needs to be moderately large to isolate oscillatory modes from wall modes in low $Pr$ fluids. This could be achieved with relatively slow rotation, or preferably, with relatively small $H$, which would additionally aid in keeping $Ra$ below $Ra_{S}$. Further, convection cells with small $H$ and large $D$ (yielding larger aspect ratio $\Gamma=D/H$) will better isolate oscillatory behavior because they mitigate effects from wall modes. The contribution from wall modes to convective heat transfer increases with increasing sidewall curvature. Supplementary Table LABEL:tab:data includes the supercriticality values for steady, wall, and oscillatory modes for all rotating cases.
Regime transitions. Thermal inertial oscillatory bulk dynamics dominate our data from just after the onset of convection (Horn and Schmid, 2017; Aurnou et al., 2018; Cheng et al., 2018; Vogt et al., 2021). A change in scaling behavior is found for sufficiently strong thermal forcing ($Ra$) near the onset of the (non-oscillatory) stationary mode of rotating convection.
Supplementary Figs. 6a, 6b, and 6c show measured $Ro_{z}$, $\theta/\Delta T$, and $Nu-1$, against the supercriticality determined by the steady mode, $Ra/Ra_{S}$. The marker notations are the same as that of Fig. 3, but vertical dotted lines are added to mark $Ra/Ra_{S}=1$. This simple demarcation works well to separate the distinctly different slope of the asymptotic regime in the data for $\theta/\Delta T$ and $Nu-1$, though it is less clear in indicating behavioral shifts in the $Ro_{z}$ data.
Supplementary Figs. 6d, 6e, and 6f show the same data against the convective Rossby number, $Ro_{c}$. The vertical dotted lines mark $Ro_{c}=1$, which has also been argued as a regime bound for rotational effects (King and Aurnou, 2013, 2015). Across all three datasets, there lacks a clear transition at a consistent value of $Ro_{c}$. There is also an aspect-ratio effect present that is instead better captured by $Ra/Ra_{S}$.
Based on Supplementary Fig. 6, we analyze the best fit scaling behaviors in our liquid metal rotating convection data using laboratory data in in the $1.5Ra_{O}^{cyl}<Ra<Ra_{S}$ range at each rotation rate ($E$) investigated.
@@ Methods
Fundamental balances. The asymptotic scaling prediction for diffusivity-free heat transfer in rotating convection is
@@
where $C_{J}\approx 1/25$ is a constant estimated in the asymptotically-reduced modeling study of Julien et al. (2012).
@@
$Re=C_{1}\left(\frac{Ra(Nu-1)}{Pr^{2}}\right)^{2/5}E^{1/5}$ (21)
(Ingersoll and Pollard, 1982; Cardin and Olson, 1994; Aubert et al., 2001; King et al., 2013; Guervilly et al., 2019; Aurnou et al., 2020; Kolhey et al., 2022; Oliver et al., 2023; Hawkins et al., 2023; Abbate and Aurnou, 2023). Further, CIA scaling arguments (Aubert et al., 2001; Abbate and Aurnou, 2023) also yield predictions for the diffusion-free rotating convective temperature perturbations in the fluid bulk, $\theta$, as
$\frac{\theta}{\Delta T}=C_{2}\left(\frac{Nu-1}{RaE}\right)^{3/5}Ro_{c}^{2/5}.$ (22)
Substituting the asymptotic heat transfer scaling (20) into the CIA predictions (21) and (22) yields estimates for the nondimensional velocity, $Re$, and the nondimensional temperature perturbation, $\theta/\Delta T$, in the asymptotic, diffusivity-free regime:
$Re=C_{1}C_{J}^{2/5}Ro_{c}^{2}E^{-1},$ (23)
@@
where $Ro=u/(2\Omega H)$ is the Rossby number. Interestingly, the asymptotic scalings (24) and (25) both depend only on the convective Rossby number, $Ro_{c}=(RaE^{2}/Pr)^{1/2}$ (cf.  Aurnou et al., 2020; Abbate and Aurnou, 2023). Setting $C_{J}\approx 1/25$ in (20), $C_{J}^{2/5}\approx(1/25)^{2/5}=0.276$ in (23), $C_{J}^{3/5}\approx(1/25)^{3/5}=0.145$ in (24), and assuming $C_{1}\approx C_{2}\approx 1$ for simplicity, we arrive at the quantitative asymptotically reduced modeling predictions for the heat transfer, momentum transfer, and bulk thermal anomalies, respectively, against which we compare our laboratory liquid metal rotating convection data in Fig. 3 and Table 2. Our results in Table 2 show that actual values are $C_{1}=1.50\pm 0.58$ and $C_{2}=2.69\pm 0.35$ for the $\Gamma\simeq 1/2$ tank, and $C_{1}=2.21\pm 0.88$ and $C_{2}=1.46\pm 0.16$ for the $\Gamma\simeq 2$ tank, which are consistent with order unity.
Earth core calculations. In this work, we have used a $\Delta T$-based framework, but the thermal flux is often what is most easily measured in most geophysical and astrophysical cases. In this case, the diffusivity-free heat transfer prediction can instead be used to predict the Rayleigh number, $Ra$, from a given flux-based Rayleigh number, $Ra_{F}$. Substituting (20) into $Ra_{F}=Ra(Nu-1)$ gives
$Ra_{F}=C_{J}Ra^{5/2}E^{2}Pr^{-1/2},$ (26)
which we can re-order as
$Ra=C_{J}^{-2/5}Ra_{F}^{2/5}E^{-4/5}Pr^{1/5}.$ (27)
The diffusivity-free velocity prediction can also be re-arranged for the Rayleigh number. Substituting the definition for the convective Rossby number, $Ro_{c}=(RaE^{2}/Pr)^{1/2}$, and re-arranging yields
$Ra=C_{J}^{-2/5}ReE^{-1}Pr.$ (28)
This provides two independent predictions for the Rayleigh number based on heat flux and velocity estimates for the geophysical or astrophysical system.
Supplementary Items
Supplementary Figure 4: Experimental apparatus. (a) Rotating convection cell. The cylindrical tank is $H$ = 40.0 cm tall by $D$ = 19.68 cm wide (aspect ratio = $\Gamma=D/H\approx 1/2$) and filled with liquid gallium ($Pr\approx 0.027$). Heat is supplied from a non-inductively wound heating pad through the bottom copper ‘heating’ block, and extracted by a heat exchanger plate in contact with the top copper ‘cooling’ block, to generate axially-aligned thermal buoyancy-driven convection. The cell is mounted on a rotating pedestal to rotate the cell about its vertical axis. An insulation blanket surrounds the cell to mitigate heat losses. (b) Underside of the top copper ‘cooling block’. Five internal thermistors of varied lengths are positioned within the fluid to measure internal temperatures. Three ultrasonic Doppler transducers are embedded vertically in the copper block to measure fluid velocities. The face of the transducer is in contact with the liquid gallium and pointed straight from the top of the cell towards the bottom.
Supplementary Figure 5: Non-rotating RBC data for each of the three cells used in the rotating study. (a) Heat transfer efficiency, $Nu$, and (b) dimensionless vertical velocity, $Re_{z}$, each versus buoyancy forcing, $Ra$. Solid lines indicate the power-law best fit to each dataset. Data for the $\Gamma\simeq 1/2$ cell is included in Supplementary Table LABEL:tab:data, data for the $\Gamma\simeq 1$ cell is included in King and Aurnou (2015), and data for the ($\Gamma\simeq 2$) cell is included in Vogt et al. (2018).
Supplementary Figure 6: Tests of transition values at which scaling behavior deviates from the diffusion-free predictions. Panels (a), (b), and (c) demonstrate a transition at approximately $Ra=Ra_{S}$ for $Ro_{z}$, $\theta/\Delta T$, and $Nu$, respectively, which is marked by a vertical dotted line. Panels (d), (e), and (f) show a lack of transition at $Ro_{c}=1$ for the same quantities, also marked by a dotted line. Solid diagonal lines in panels (c), (d), and (e) indicate the diffusivity-free predictions.
Supplementary Table 3: Liquid gallium data for the $\boldsymbol{\Gamma\simeq 1/2}$ cell.

### v2

@@ Abstract
Table 2: Power law fits corresponding to the data shown in Fig. 3. Prefactors ($A$, $B$, $C$) and exponents ($\alpha$, $\beta$, $\gamma$) of unity correspond to the result of the plane layer asymptotically reduced models of [10].
@@
Fig. 3 compares our measurements with the theoretical DF predictions formulated with Julien et al. [10]’s $C_{J}=1/25$ prefactor. The data is presented as colored markers, with a solid black line indicating the asymptotically reduced predictions. Power law fits to the oscillatory data (solid points) are shown as dotted lines of the same color. Inset plots display the best-fit values, and Table 2 lists the specific best-fit power law relations for each dataset. Pre-factors and exponents of unity, which are marked by a solid black horizontal line in the inset plots, correspond to the asymptotically reduced result. The oscillatory data is in good quantitative agreement with this prediction. Hollow points (indicating $Ra>Ra_{S}$) clearly deviate from the diffusion-free scaling. Thus, the thermal-inertial modes follow the non-diffusive scalings, whilst the inclusion of low to moderate supercriticality bulk stationary modes truncates this behavior.
The heat transfer efficiency ($Nu$) results in Fig. 3a show a close fit between the asymptotically reduced and $\Gamma\simeq 2$ cell data. Decreasing the aspect ratio likely introduces strong wall mode effects that enhance global heat transfer efficiency [23, 24, 25, 26], causing the slope to deviate from the asymptotic scaling. The inset plot, which shows the best-fit values, reveals a clear trend toward the asymptotic limit as the aspect ratio increases for both the scaling exponent and the prefactor. Thus, the wider aspect ratio cell more closely replicates infinite plane layer behavior than the narrower aspect ratio cells, making it a better fit to the asymptotically reduced model result.
The thermal anomaly ($\theta$) results shown in Fig. 3b exhibit excellent agreement with theory across both aspect ratios ($\Gamma\simeq 1/2$, $\Gamma\simeq 2$). The inset plots show scaling exponents near unity for both datasets. Unlike global heat transfer, internal fluctuations appear to be insensitive to wall flow behaviors such that they exclusively follow the thermal-inertial bulk scaling trends irrespective of aspect ratio value.
The vertical velocity $Re_{z}$ results shown in Fig. 3c also strongly align with DF theory, as they are similarly unaffected by wall modes. Overall, we find simultaneous agreement with diffusion-free predictions across all three independently measured quantities in this study: thermal anomalies in the bulk, velocities in the bulk, and global heat transfer. Further, we find increasing fidelity of this agreement in higher aspect ratio containers [cf. 40].
These experiments show that diffusion-free physics can exist even in relatively small, closed containers of rotating fluid. Of specific note, we find that rotating low $Pr$ liquid metal convection in a relatively short, large aspect ratio tank ($\Gamma\simeq 2$) robustly agrees with the varied $Pr$ results of the asymptotically-reduced planar modeling study of Julien et al. [10]. This work thus verifies the $C_{J}$ constant found by Julien et al. [10] to hold for DF planar systems with no-slip top and bottom boundaries. In our laboratory setup, we utilize the oscillatory regime of low $Pr$ rotating convection to achieve this goal. However, Julien et al. [10] find that the scaling can hold for a range of $Pr$ assuming the criteria of rotationally-constrained turbulence are met ($Ro\ll 1\ll Re$), as expected in planetary interior flows [41, 42] and in many stellar settings as well [43]. Further, some numerical studies conducted with different geometries and values of $Pr$ have also found agreement with diffusion-free scalings, albeit with differing scaling coefficients [e.g., 44, 45, 46].
@@
Measurements of the inertial oscillation-dominated range of liquid metal RRBC in closed cylindrical laboratory containers reveal that asymptotic DF heat and momentum transfer can be realized on the laboratory scale. We observe these trends in measurements of internal flow velocities and thermal fluctuations regardless of tank geometry. In contrast, global heat transfer is significantly influenced by wall modes in narrow, low aspect ratio cells, with asymptotic trends predominantly found in wide, high aspect ratio cells that better resemble plane layer geometry. If wall modes can be suppressed [cf. 47], it may be possible to achieve asymptotic heat transfer scalings in taller, more slender $\Gamma\lesssim 1$ tanks. It may be possible to further investigate DF scaling behaviors using lower $Pr$ fluids, such as liquid sodium, in which $Ra_{S}/Ra_{O}$ attain larger values. This work verifies the existence of diffusivity-free convection in a standard desktop-scale RRBC experiment and demonstrates its potential as a laboratory analog for investigating the buoyancy-driven turbulence that shapes the convection zone dynamics of planets and stars.
J.A.A. acknowledges graduate funding from the DoD’s NDSEG Fellowship program. J.M.A. thanks the NSF Geophysics Program for support via award #2143939.
Supplementary Material
The supplement consists of A) an explanation of relevant scaling equations for diffusivity-free rotating convection and B) a data table containing the new experimental data used in this study.
.1 Fundamental Scalings of Diffusivity-Free Rotating Convection
Heat Transfer.
@@
The diffusivity-free (DF) scaling prediction for superadiabatic heat transfer in rotating convection follows
$Nu_{conv}\sim\ Ra^{3/2}E^{2}Pr^{-1/2}$ (4)
[cf. 48, 10].The subscript ‘conv’ denotes the superadiabatic, or ‘convective’, quantity such that
$Nu_{conv}=\frac{q_{conv}}{q_{cond}}\ ,$ (5)
where $q_{conv}$ and $q_{cond}$ represent the convective heat flux and heat lost to conduction, respectively. These quantities are related to the total Nusselt number, $Nu$, following
$Nu=\frac{q_{total}}{q_{cond}}=\frac{q_{cond}+q_{conv}}{q_{cond}}=1+\frac{q_{conv}}{q_{cond}}\ ,$ (6)
implying that $Nu-1=Nu_{conv}$. In systems where $Nu\gg 1$, it is usually assumed that $Nu-1\approx Nu$ and therefore $Nu_{conv}\approx Nu$, but this assumption is not valid near the onset of convection when $Nu$ is small, as is the case in this study. Substituting $Nu$ into (4) and including a prefactor yields the form of the DF prediction used to analyze our results:
@@
where $C_{J}\approx 1/25$ is the constant estimated in the asymptotically-reduced modeling study of Julien et al. [10].
This formulation can be further explored by considering the heat flux contributions. The convective and conductive components of the total flux can be mathematically defined following
$q_{conv}=\rho C_{p}\langle u_{z}\theta\rangle\quad\text{and}\quad q_{cond}=k\Delta T/H\ ,$ (8)
where $\rho$ is the fluid density, $C_{p}$ is the specific heat capacity, $k$ is the thermal conductivity, $\Delta T$ is the global temperature differential, $H$ is the fluid layer height, $u_{z}$ is the vertical flow velocity, $\theta$ is the fluid temperature perturbation, and delimited-⟨$\langle$..$\rangle$ … ⟩ indicates a turbulent time and volume average. Substituting (8) into (6) and re-arranging yields
$Nu-1=\frac{q_{conv}}{q_{cond}}=\frac{\rho C_{p}\langle u_{z}\theta\rangle}{k\Delta T/H}\sim RePr\frac{\theta}{\Delta T}\ .$ (9)
Equation (9) demonstrates that the convective heat transfer efficiency can be estimated by $Nu_{conv}\sim RePr(\theta/\Delta T)$ in addition to $Nu_{conv}\sim Nu-1$. Supplementary Fig. 4 shows each of these estimates against the DF scaling prediction from (4) for the $\Gamma=1/2$ and $\Gamma=2$ convection cells. In this study, we independently measure the global heat flux ($q_{total}$), flow velocity ($u_{z}$), and temperature fluctuations ($\theta$), ensuring that the data represented in each panel of Supplementary Fig. 4 is uniquely acquired. Panel (a) plots the superadiabatic heat transfer determined from the global heat flux measurements, as was done in Fig. 3a. The wide $\Gamma=2$ cell maintains excellent agreement, while the narrower $\Gamma=1/2$ cell deviates due to the extensive contribution of wall modes to the global convective heat transfer. Panel (b) plots the superadiabatic heat transfer estimated from our internal measurements of flow velocity and temperature fluctuations. These interior quantities are indifferent to wall modes, thus data from both aspect ratio cells align with the DF heat transfer scaling in the oscillatory regime. It is thus important to consider the convective heat transfer, rather than the total heat transfer, in these low $Nu$ experiments.
SUPP. FIG. 4: Superadiabatic $Nu_{conv}$ determined from measurements of (a) global heat flux ($q_{total}$) or (b) internal velocity ($u_{z}$) and temperature fluctuations ($\theta$), plotted against the diffusivity-free scaling prediction given by (4).
Internal Quantities.
Supplementary Fig. 4b considers the internal quantities ($u_{z}$ and $\theta$) in relation to the heat transfer efficiency scaling of (4). However, in Fig. 3, we separately show the flow velocities (Fig. 3b) and the temperature fluctuations (Fig. 3c) against their independent scaling predictions.
@@
$Re=C_{1}\left(\frac{Ra(Nu-1)}{Pr^{2}}\right)^{2/5}E^{1/5}\ ,$ (10)
where $C_{1}$ is an arbitrary constant [11, 49, 50, 51, 44, 52, 53, 17, 54, 9]. Further, CIA scaling arguments also yield predictions for the diffusion-free rotating convective temperature perturbations in the fluid bulk, $\theta$, as
$\frac{\theta}{\Delta T}=C_{2}\left(\frac{Nu-1}{RaE}\right)^{3/5}Ro_{c}^{2/5}\ ,$ (11)
where $C_{2}$ is also an arbitrary constant [9]. Substituting the asymptotic heat transfer scaling (7) into the CIA predictions, (10) and (11), yields estimates for the nondimensional velocity, $Re$, and the nondimensional temperature perturbation, $\theta/\Delta T$, in the asymptotic, diffusivity-free regime:
$\vspace{-5mm}Re=C_{1}C_{J}^{2/5}Ro_{c}^{2}E^{-1}\ ,$ (12)
@@
where $Ro=u/(2\Omega H)$ is the Rossby number. Interestingly, the asymptotic scalings (13) and (14) both depend only on the convective Rossby number, $Ro_{c}=(RaE^{2}/Pr)^{1/2}$ [cf.  52, 9]. Setting $C_{J}\approx 1/25$ in (7), $C_{J}^{2/5}\approx(1/25)^{2/5}=0.276$ in (12), $C_{J}^{3/5}\approx(1/25)^{3/5}=0.145$ in (13), and assuming $C_{1}\approx C_{2}\approx 1$ for simplicity, we arrive at the quantitative asymptotically reduced modeling predictions for the heat transfer, momentum transfer, and bulk thermal anomalies, respectively, against which we compare our laboratory liquid metal rotating convection data in Fig. 3 and Table 2. Our results in Table 2 show that actual values are $C_{1}=1.50\pm 0.58$ and $C_{2}=2.69\pm 0.35$ for the $\Gamma\simeq 1/2$ tank, and $C_{1}=2.21\pm 0.88$ and $C_{2}=1.46\pm 0.16$ for the $\Gamma\simeq 2$ tank, which are consistent with order unity.
.2 Data Table
SUPP. TABLE 3: Liquid gallium data for the $\Gamma\simeq 1/2$ cell.